OpenStudy (yamyam70):

i need help with probabilities A foreign student club lists as its members 2 Canadians, 3 Japanese, 5 Italians, and 2 Germans. If a committee of 4 is selected at random, find the probability that all nationalities except the Italians are represented

OpenStudy (anonymous):

The total number of students is 12, so to find the probability we'll do the following for the Canadians the probability is 2/12 for the next Japanese member 3/11 for the german member 2/10 and for the last member 4/9 as there are only 4 non italian members left multiplying the probabilities 2/12*3/11*2/10*4/9=0,00404=0,404%

OpenStudy (nubeer):

P(all nationalities except the Italians are represented) = 7C4/12C4

OpenStudy (kropot72):

Canadians Japanese Germans Italians (A) 1 1 2 0 (B) 1 2 1 0 (C) 2 1 1 0 The possible combinations are shown above. \[\large P(A)=\frac{2C1\times3C1\times2C2\times5C0}{12C4}=\frac{6}{495}\] \[\large P(B)=\frac{2C1\times3C2\times2C1\times5C0}{12C4}=\frac{12}{495}\] \[\large P(C)=\frac{2C2\times3C1\times2C1\times5C0}{12C4}=\frac{6}{495}\] The events A, B and C are mutually exclusive. Therefore the probability that all nationalities except the Italians are represented is: \[\large P(A)+P(B)+P(C)=\frac{6}{495}+\frac{12}{495}+\frac{6}{495}=\frac{24}{495}=0.0485\]

OpenStudy (yamyam70):

Thank you so much, I passed my statistics class because of this, sorry for the late credit :))

OpenStudy (kropot72):

That you for the feedback. Congratulations for your result :)